# supplementary

### v2

@@ 7.2.2 Investment Demand
The iterative algorithm of Zellner (1962) takes 52 iterations to converge to a local mode, with a sum of squared residuals (SSR) of 216,943 and profile log-likelihood of -263.7. We run ALPS with seven inverse-temperature levels (1.00, 1.10, 1.40, 1.96, 2.74, 3.84, 5.38) and hot state 1/15 = 0.067. Three modes are located as shown in Figure 5 (d), although the third mode has negligible probability mass. The first mode discovered by the EC part of ALPS matches the estimate from ‘systemfit,’ while the other two modes have profile log-likelihoods of -264.9 and -329.1, respectively.
@@
This problem is particularly ill-posed due to parameter unidentifiability (which leads to the multimodality in the posterior). In this problem, despite attempts to reparameterize, there remain issues with the modes of the posterior distribution being located on long, thin ridges that seem to decay with significantly heavier tails than Gaussian. This means that to obtain a good Laplace approximation to the mode the ALPS procedure would need to use very large inverse-temperature values. However, when this was attempted it led to severe computational issues with regards to machine precision, likely due to the condition numbers of the Hessians at the mode points, illustrating the ill-posed nature of the problem. To overcome this a truncated version of the annealed HAT targets was utilised for annealed temperatures with $\beta>1$. These are a small modification to the HAT targets and they are defined and discussed in detail in Appendix LABEL:sec:Truncated.
@@ 7.3 Example 3: Spectral Density Model
We used two inverse-temperature levels (1.0, 1.5) and hot state 0.025. The exploration component ALPS-EC took 12513 iterations to find 50 modes, as shown in Fig. 6 (b), although only 7 of these modes have non-negligible probability mass. It took 16.5 minutes for 80,000 iterations of ALPS.
@@ 8 Conclusion and Further Work
There are a number of further considerations regarding adding robustness and practicality to the ALPS procedure. Many of the ideas for exploration are already alluded to in the paper: improving the effectiveness of the EC part by running more than one exploration chain and having multiple temperature levels; parallel implementation of the PA and EC parts of ALPS; and introducing weightings to regions to stabilise the inter-modal masses when the HAT targets don’t sufficiently preserve the weights. Adapting the ideas of ALPS to facilitate an automatic and scalable reversible jump MCMC algorithm is the immediate focus for further work.

### v3

@@ 7.2.2 Investment Demand
The iterative algorithm of Zellner (1962) takes 52 iterations to converge to a local mode, with a sum of squared residuals (SSR) of 216,943 and profile log-likelihood of -263.7. We run ALPS with $\beta_{\text{max}}=5.38$ and seven inverse-temperature levels (1.00, 1.10, 1.40, 1.96, 2.74, 3.84, 5.38) and hot state 1/15 = 0.067. Three modes are located as shown in Figure 6 (d), although the third mode has negligible probability mass. The first mode discovered by the EC part of ALPS matches the estimate from ‘systemfit,’ while the other two modes have profile log-likelihoods of -264.9 and -329.1, respectively.
@@
This problem is particularly ill-posed due to parameter unidentifiability (which leads to the multimodality in the posterior). In this problem, despite attempts to reparameterize, there remain issues with the modes of the posterior distribution being located on long, thin ridges that seem to decay with significantly heavier tails than Gaussian. This means that to obtain a good Laplace approximation to the mode the ALPS procedure would need to use very large inverse-temperature values. However, when this was attempted it led to severe computational issues with regards to machine precision, likely due to the condition numbers of the Hessians at the mode points, illustrating the ill-posed nature of the problem. To overcome this a truncated version of the annealed HAT targets was utilised for annealed temperatures with $\beta>1$. These are a small modification to the HAT targets and they are defined and discussed in detail in Appendix B.2.
@@ 7.3 Example 3: Spectral Density Model
We used two inverse-temperature levels with the target inverse temperature 1.0 and $\beta_{\text{max}}=1.5$ as well as a hot state 0.025. The exploration component ALPS-EC took 12513 iterations to find 50 modes, as shown in Fig. 7 (b), although only 7 of these modes have non-negligible probability mass. It took 16.5 minutes for 80,000 iterations of ALPS.
@@ 8 Conclusion and Further Work
There are a number of further considerations regarding adding robustness and practicality to the ALPS procedure. Many of the ideas for exploration are already alluded to in the paper: improving the effectiveness of the EC part by running more than one exploration chain and having multiple temperature levels; parallel implementation of the PA and EC parts of ALPS; and introducing weightings to regions to stabilise the inter-modal masses when the HAT targets don’t sufficiently preserve the weights. Another possibility would be to incorporate the cold state level into a simulated tempering framework, interpolating between the cold Laplace approximation and the target distribution, and serving, per se, as a prior or reference distribution as in Surjanovic et al. (2022). This would allow perfect regeneration of the Markov chain at each visit to the cold level thanks to the i.i.d. sampling of the Laplace approximation. Adapting the ideas of ALPS to facilitate an automatic and scalable reversible jump MCMC algorithm is the immediate focus for further work.
Acknowledgements